# Infrequent Resolving Algorithm for Online Linear Programming

### v3

@@ 1.1 Literature Review
Known distribution. We now review the stream assuming the knowledge of the arrival probabilities at the beginning, which is widely studied in the network revenue management (NRM) problem. Originating from the airline industry, NRM has garnered significant attention from both academia and industry. There are two mainstreams of research, price-based NRM and quantity-based NRM. The former one studies the dynamic pricing problem under resource constraints (see Gallego and Van Ryzin 1994); the latter one studies the dynamic resource allocation problem (see Talluri and Van Ryzin 1998). Our work is closely related to the quantity-based NRM, in which most algorithms are LP-based. For example, Talluri and Van Ryzin (1998) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the bid-price control (BPC) policy, which uses the optimal dual variable to the fluid model as the values of resources and accepts a request if and only if the offered price is larger than the total value of resources. Cooper (2002) proves an $\mathcal{O}(\sqrt{T})$ regret bound for the booking-limit control (BLC) policy, which assigns quotas to each request according to the fluid model and accepts a request until the corresponding quota is depleted. Reiman and Wang (2008) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the probabilistic allocation control (PAC) policy, which probabilistically accepts a request according to the ratio of the fluid model solution to the expected demand.
@@
Table 2: Comparison among Algorithms for Known-Probability Case.
Paper
Regret
(Non-degenerate)
Regret
# of Resolvings
@@
Jasin and Kumar (2012) $\mathcal{O}(1)$ [^3] Bumpensanti and Wang (2020) prove that the regret of the algorithm in Jasin and Kumar (2012) is $\Theta(\sqrt{T})$. $\mathcal{O}(\log T)$
Bumpensanti and Wang (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
Arlotto and Gurvich (2019)
(Single constraint )
$\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Vera and Banerjee (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Vera et al. (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Sun et al. (2020) $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ 0
Reiman and Wang (2008) $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ 2
Jasin and Kumar (2012) $\mathcal{O}(T^{\frac{1}{2^{M}}})$ $M$
Bumpensanti and Wang (2020) $\mathcal{O}(T^{5/12})$ $\mathcal{O}(T^{5/12})$ 2
Gupta (2024) $\mathcal{O}(1)$ 1
This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $M$
@@
Observing the potential to resolve the fluid model to reduce the regret, many works consider algorithms with resolving. For example, Cooper (2002) provides an example where frequent resolving may hurt the revenue. Then, considering the asymptotic regime, Reiman and Wang (2008) prove that a single resolving can reduce the regret of the PAC policy from $\Theta(\sqrt{T})$ to $o(\sqrt{T})$. Jasin and Kumar (2012) and Jasin and Kumar (2013) show that the PAC policy can significantly benefit from resolving, reducing its regret from $\Theta(\sqrt{T})$ to $\mathcal{O}(1)$, while neither BPC nor BLC can benefit. Specifically, under the non-degeneracy assumption, the PAC policy with periodic ($\mathcal{O}(T)$ times) or midpoint ($\mathcal{O}(\log T)$ times) resolving can achieve $\mathcal{O}(1)$ regret. Without the non-degeneracy assumption, Bumpensanti and Wang (2020) show that the above regret in general is $\Omega(\sqrt{T})$, and provide a modified PAC policy with infrequent (i.e., $\mathcal{O}(\log\log T)$ times) resolving that guarantees $\mathcal{O}(1)$ regret. Recently, Arlotto and Gurvich (2019) study an alternative interpretation of the fluid solution, which accepts the request if and only if the ratio of the primal solution to the expected demand is no less than $1/2$. They prove an $\mathcal{O}(1)$ regret bound for the multi-secretary problem. In our work, we refer to such policy as the “argmax policy” because it takes the action (accept/reject) with the larger value in the primal solution. Vera and Banerjee (2021) and Vera et al. (2021) generalize this idea to the multi-constraint problem, but their policy requires solving LPs in every period. In our work, we also use the argmax policy as the base policy. We show that such a policy can guarantee $\mathcal{O}(1)$ regret by solving only $\mathcal{O}(\log\log T)$ LPs. In the LP-free regime, Sun et al. (2020) propose an LP-free algorithm with $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ regret.
@@
Finally, we would like to mention some results on the case when the decision maker is only allowed to solve a finite number (i.e., $M$) of LPs under the NRM model. For this problem, Reiman and Wang (2008) show that the regret can be reduced to $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ if we can solve LPs twice. However, the proof cannot be directly extended to the multiple-resolving case. Under the non-degeneracy assumption, Jasin and Kumar (2012) prove that solving LP at $M$ periods can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound. Due to the restricted choice of resolving times, Bumpensanti and Wang (2020) prove that the regret bound of their policy is $\mathcal{O}(T^{5/12})$ given $M=2$, which is worse than previous results. In addition, Gupta (2024) recently proposes an algorithm which only solves LP once at the beginning and then greedily makes decisions to minimize a sum-of-squares function in each period, whose regret is proved to be $\mathcal{O}(1)$ under the non-degeneracy assumption. Compared to the above literature, our results can meet the single-resolving result of Reiman and Wang (2008) and extend it to the finite-resolving case. We derive similar regret bounds in Jasin and Kumar (2012) but allow for degenerate cases. A summary of the existing results and our result is presented in Table 2 (where $\epsilon$ can be any positive constant).
@@ 2.1 Argmax with Infrequent Resolving (AIR) Policy
We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
@@
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
We now explain the intuition of Algorithm 1. The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed. At each time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ as an approximation to the future demand and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$. The optimal solution $\boldsymbol{y}^{*}$ of the fluid LP provides an approximation to the optimal policy. At each time period $t\notin\mathcal{T}$, the decision maker updates an approximation $\boldsymbol{u}^{t}$ to the optimal solution in a “greedy” fashion as described in the algorithm. We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions. At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand. Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected. In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@ 2.2 Resolving Schedule
As Figure 1 shows, the learning time set $\mathcal{T}_{L}$ is concentrated at the beginning of the time horizon. This is similar to the approach in Agrawal et al. (2014) and Gupta and Molinaro (2014), but our method requires solving fewer LPs, i.e., $\mathcal{O}(\log\log T)$ times compared to $\mathcal{O}(\log T)$ times in Agrawal et al. (2014) and Gupta and Molinaro (2014). Since empirical estimations fluctuate drastically at the beginning of the time horizon, the learning time set is designed to update the estimations promptly to avoid the accumulation of error. In contrast to the learning time set, the approximation time set $\mathcal{T}_{A}$ is concentrated at the end of the time horizon. It shares some similarities with the NRM literature, e.g., Jasin and Kumar (2012) and Bumpensanti and Wang (2020). Since the solution to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ usually changes drastically when the inventory is scarce, it is crucial to have a good approximation of the optimal solution at the end of the time horizon. The approximation time set is thus designed to control the deviation of the approximated solution from the true optimal fluid solution. The total number of resolving is $|\mathcal{T}|=K_{L}+K_{A}+1=\mathcal{O}(\log\log T)$. Then, we present our main result in the following theorem.
@@
According to Theorem 1, the AIR algorithm can achieve a constant regret by solving $\mathcal{O}(\log\log T)$ LPs. In the following, we present the proof idea of Theorem 1.
@@ 2.3 Proof of Regret Bound
According to Proposition 2, the demand estimation error is relatively small with high probability. Both bounds consist of two parts: The first term is due to the gap between the empirical probability $\hat{\boldsymbol{p}}^{t}$ and the underlying true probability $\boldsymbol{p}$; the second term is due to the deviation of the random future demand $\boldsymbol{Z}^{t}$ from the expected demand $(T-t+1)\hat{\boldsymbol{p}}^{t}$.
Step 3: Surrogate LP for $\boldsymbol{u}^{t}$. In the literature on argmax policies (e.g., Vera and Banerjee 2021 and Vera et al. 2021), the proposed policy makes decisions based on the optimal solution of the fluid problem at each period. This approach allows them to bridge the proposed policy with the optimal policy using two LPs: the fluid problem and the hindsight problem. However, we solve the fluid problem only at a few selected periods, and approximate the optimal solution by $\boldsymbol{u}^{t}$ in other periods. In this case, $\boldsymbol{u}^{t}$ is not an optimal solution to the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ for most time periods. Consequently, the linkage between the fluid model and the proposed policy breaks down, and the proof technique in the literature cannot directly apply to our proof. To overcome this challenge, we prove that $\boldsymbol{u}^{t}$ is a good approximation to the optimal solutions of the fluid problem by introducing a surrogate LP. In the following, we show that $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP with high probability.
@@
As Proposition 3 shows, the approximated solution $\boldsymbol{u}^{t}$ is optimal to a surrogate LP, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, with high probability. Recall that we prove that the demand estimation $\boldsymbol{d}^{t}$ is not far away from the random demand $\boldsymbol{Z}^{t}$. Then, we will prove that the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ is not far away from the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$. Thus, we can show that if $u_{j}^{t}\geq d_{j}^{t}/2$, then there exists an optimal solution $\boldsymbol{y}^{*}$ of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$ such that $y^{*}_{j}\geq 1$ with high probability, implying that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j,1)=0$ according to Proposition 1. The logic for the case with $u_{j}^{t}<d_{j}^{t}/2$ is similar.
@@
According to Proposition 4, for most periods, under the good event, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j^{t}}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$, resulting in $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$. Therefore, the probability $\mathbb{P}\left(\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})>0\right)$ is upper bounded by the bad event probability. Finally, with the above four steps, we are prepared to prove Theorem 1 by bounding the right-hand side of (5).

### v4

@@ 1.1 Literature Review
Known distribution. We now review the stream assuming the knowledge of the arrival probabilities at the beginning, which is widely studied in the network revenue management (NRM) problem. Originating from the airline industry, NRM has garnered significant attention from both academia and industry. There are two mainstreams of research, price-based NRM and quantity-based NRM. The former one studies the dynamic pricing problem under resource constraints (see Gallego and Van Ryzin 1994); the latter one studies the dynamic resource allocation problem (see Talluri and Van Ryzin 1998). Our work is closely related to the quantity-based NRM, in which most works adopt the finite-support assumption (see Jiang et al. 2022 and Besbes et al. 2024 for some exceptions). Most early works only solve the fluid LP at the beginning and then design a policy based on the optimal solution. For example, Talluri and Van Ryzin (1998) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the bid-price control (BPC) policy, which uses the optimal dual variable to the fluid model as the values of resources and accepts a request if and only if the offered price is larger than the total value of resources. Cooper (2002) proves an $\mathcal{O}(\sqrt{T})$ regret bound for the booking-limit control (BLC) policy, which assigns quotas to each request according to the fluid model and accepts a request until the corresponding quota is depleted. Reiman and Wang (2008) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the probabilistic allocation control (PAC) policy, which probabilistically accepts a request according to the ratio of the fluid model solution to the expected demand.
@@
Table 2: Comparison among Algorithms for Known-Probability Case (Finite-Support Distribution).
@@
Paper Regret (Non-degenerate) Regret (Degenerate) # of Resolvings Jasin and Kumar (2012) $\mathcal{O}_{\delta}(1)$ [^3] Bumpensanti and Wang (2020) prove that the regret of the algorithm in Jasin and Kumar (2012) is $\Theta(\sqrt{T})$. $\mathcal{O}(\log T)$ Bumpensanti and Wang (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ Vera and Banerjee (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Vera et al. (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Banerjee and Freund (2024) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ (In Expectation) Reiman and Wang (2008) $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ 2 Jasin and Kumar (2012) $\mathcal{O}_{\delta}(T^{\frac{1}{2^{M}}})$ $M$ Bumpensanti and Wang (2020) $\mathcal{O}(T^{5/12})$ $\mathcal{O}(T^{5/12})$ 2 Sun et al. (2020) $\mathcal{O}_{\delta}(T^{3/8}(\log T)^{5/4})$ 0 Gupta (2024) $\mathcal{O}_{\delta}(1)$ 1 This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $M$
@@
Observing the potential to resolve the fluid model to reduce the regret, many works consider LP-based algorithms with more resolvings. For example, Reiman and Wang (2008) prove that a single resolving can reduce the asymptotic regret of the PAC policy from $\Theta(\sqrt{T})$ to $o(\sqrt{T})$. Jasin and Kumar (2012) and Jasin and Kumar (2013) show that the PAC policy can significantly benefit from resolving, while neither BPC nor BLC can benefit. Specifically, under the non-degeneracy assumption, the PAC policy with periodic ($\mathcal{O}(T)$ times) or midpoint ($\mathcal{O}(\log T)$ times) resolving can achieve $\mathcal{O}_{\delta}(1)$ regret. Without the non-degeneracy assumption, Bumpensanti and Wang (2020) show that the above regret in general is $\Omega(\sqrt{T})$, and provide a modified PAC policy with infrequent (i.e., $\mathcal{O}(\log\log T)$ times) resolving that guarantees $\mathcal{O}(1)$ regret. Note that the resolving schedule in Bumpensanti and Wang (2020) shares some similarities with the second half of our schedule, but their proof framework cannot deal with the case without distribution knowledge. Recently, Arlotto and Gurvich (2019) study an alternative interpretation of the fluid solution, which accepts the request if and only if the ratio of the primal solution to the expected demand is no less than $1/2$. They prove an $\mathcal{O}(1)$ regret bound for the multi-secretary problem. In our work, we refer to such policy as the “argmax policy” because it takes the action (accept/reject) with the larger value in the primal solution. Then, Vera and Banerjee (2020) and Vera et al. (2021) generalize this idea to the multi-constraint problem, but their policy requires solving LPs in every period. Subsequently, Banerjee and Freund (2024) propose a constant-regret algorithm whose expected number of resolvings is $\mathcal{O}(\log\log T)$. Similarly, the proof in Banerjee and Freund (2024) cannot handle the unknown-probability case under infrequent resolving. In our work, inspired by the argmax policy, we propose a constant-regret policy whose resolving schedule can be determined at the beginning and the number of resolvings is $\mathcal{O}(\log\log T)$.
@@
Then, we would like to mention some results on the case when the decision maker is only allowed to solve a finite number (i.e., $M$) of LPs under the NRM model. For this problem, Reiman and Wang (2008) show that the regret can be reduced to $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ if we can solve LPs twice. However, the proof cannot be directly extended to the multiple-resolving case. Under the non-degeneracy assumption, Jasin and Kumar (2012) prove that solving LP at $M$ periods can induce an $\mathcal{O}_{\delta}(T^{\frac{1}{2^{M}}})$ regret bound. Due to the restricted choice of resolving times, Bumpensanti and Wang (2020) prove that the regret bound of their policy is $\mathcal{O}(T^{5/12})$ given $M=2$, which is worse than previous results. Sun et al. (2020) propose an LP-free algorithm with $\mathcal{O}_{\delta}(T^{3/8}(\log T)^{5/4})$ regret under the non-degeneracy assumption. In addition, Gupta (2024) recently proposes an algorithm which only solves LP once at the beginning (to find the optimal basis) and then greedily makes decisions to minimize some function in each period, whose regret is proved to be $\mathcal{O}_{\delta}(1)$ under the non-degeneracy assumption. Compared to the above literature, our results can meet the single-resolving result of Reiman and Wang (2008) and extend it to the finite-resolving case. We derive similar regret bounds in Jasin and Kumar (2012) but allow for degenerate cases. A summary of the existing results and our result is presented in Table 2 (where $\epsilon$ can be any positive constant).
@@ 2.1 Argmax with Infrequent Resolving (AIR) Policy
We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry, and $[n]$ to denote the set $\{1,2,3,\dots,n\}$. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
@@
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$, $u_{j}^{t}>1$, and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
We now explain the intuition of Algorithm 1. The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed. At each resolving time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$. To implement our policy, we need to maintain two approximations in our algorithm: The vector $\boldsymbol{u}^{t}=\boldsymbol{y}^{*}$ approximates the numbers of accepted future customers (of different types) under the optimal policy, and the vector $\boldsymbol{d}^{t}=(T-t+1)\hat{\boldsymbol{p}}$ approximates the numbers of future customer arrivals (of different types). However, at each non-resolving time period $t\notin\mathcal{T}$, the decision maker cannot access the optimal solution $\boldsymbol{y}^{*}$ for the current period. In this case, we will update the approximations $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$ by simple first-order computations in a “greedy” fashion as described in the algorithm: Once a type-$j$ customer arrives at period $t-1$, we subtract the number of future type-$j$ arrivals by one, i.e., $\boldsymbol{d}^{t}=\boldsymbol{d}^{t-1}-\boldsymbol{e}_{j}$; Once a type-$j$ customer is accepted at period $t-1$, we subtract the number of future accepted type-$j$ arrivals by one, i.e., $\boldsymbol{u}^{t}=\boldsymbol{u}^{t-1}-\boldsymbol{e}_{j}$.
Given the approximations $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$, we adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera et al. (2021) to make accept/reject decisions. At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$. Note that $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand. Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if and only if more should be accepted than rejected (i.e., $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$). In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@ 2.2 Resolving Schedule
As Figure 1 shows, the learning time set $\mathcal{T}_{L}$ is concentrated at the beginning of the time horizon. This is similar to the approach in Agrawal et al. (2014) and Gupta and Molinaro (2014), but our method requires solving fewer LPs, i.e., $\mathcal{O}(\log\log T)$ times compared to $\mathcal{O}(\log T)$ times in Agrawal et al. (2014) and Gupta and Molinaro (2014). Since empirical estimations fluctuate drastically at the beginning of the time horizon, the learning time set is designed to update the estimations promptly to avoid the accumulation of learning error. In contrast to the learning time set, the approximation time set $\mathcal{T}_{A}$ is concentrated at the end of the time horizon. It shares some similarities with the NRM literature, e.g., Jasin and Kumar (2012) and Bumpensanti and Wang (2020). At the end of the time horizon, the solution to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ usually changes drastically because of the scarce inventory. Since the first-order computations may incur significant errors in this case, the approximation time set is thus designed to control the deviation of the approximated solution from the true optimal fluid solution. The total number of resolving is $|\mathcal{T}|=K_{L}+K_{A}+1=\mathcal{O}(\log\log T)$. Then, we present our main result in the following theorem.
@@
According to Theorem 1, the AIR algorithm can achieve a constant regret by solving $\mathcal{O}(\log\log T)$ LPs. Before proceeding, we provide an overview of the proof idea of Theorem 1 in four steps. First, we decompose the regret by periods, and identify events under which a revenue loss occurs. By doing so, we reduce the analysis to bounding the probability of such events. Second, we prove that $\boldsymbol{d}^{t}$ in Algorithm 1 is a good approximation of future arrival numbers. Third, we identify that $\boldsymbol{u}^{t}$ is an optimal solution of a surrogate LP with high probability, and hence can show that $\boldsymbol{u}^{t}$ is a good approximation of the optimal solution of the fluid problem. Lastly, given that $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$ are good approximations, we can prove that the event probability in the first step is small under the argmax decision, resulting in a constant regret. In the following subsection, we present a more detailed proof of Theorem 1 step by step. Before that, we first provide remarks to compare our techniques with the literature reviewed in Section 1.1 to highlight our contribution.
Remark 1 (Comparison with Literature).
First, Bumpensanti and Wang (2020) consider the problem where the probabilities are known at the beginning, and provide a constant-regret algorithm with an $\mathcal{O}(\log\log T)$ resolving schedule similar to our approximation time set. They use the thresholding technique to adjust the acceptance probabilities in the PAC control: When the probability is lower (higher, resp.) than a threshold, the probability will be adjusted to 0 (1, resp.). The high-level philosophy of the thresholding policy is similar to the argmax policy, which can be seen as a thresholding action with $0.5$ as the threshold. However, different from our policy, their policy between two resolving periods is static (i.e., the acceptance probabilities are not updated), and the induction proof in Bumpensanti and Wang (2020) highly relies on the knowledge of arrival probabilities. Even if we complement the resolving time set with our learning time set, their proof cannot easily be adapted to the unknown-probability case. Indeed, in Bumpensanti and Wang (2020), they propose a way to deal with the unknown-probability case by requiring per-period resolving rather than infrequent resolving.
Second, for the known-probability case, Banerjee and Freund (2024) provide a lazy-resolving algorithm achieving a constant regret with $\mathcal{O}(\log\log T)$ expected number of resolvings. Their algorithm sets lower confidence bounds for the numbers of accepted customers of different types based on the optimal solution of the fluid model, and updates the bounds when the number of accepted customers exceeds the corresponding bound. The proof cannot be extended to the unknown-probability case because the lower confidence bounds is sensitive to the estimation error of the arrival probabilities.
Third, for the unknown-probability case, Xie et al. (2023) achieve constant regret by implementing the argmax policy with per-period resolving. However, the proof cannot be directly used for our infrequent resolving schedule because the empirical estimation is infrequently updated and hence the optimal solution to the fluid model is not available in most periods. To deal with this problem, we provide a novel idea to identify a surrogate LP which admits the approximated solution as an optimal solution with high probability, such that the approximated solution can be compared with the optimal solution of the fluid model. ∎
@@ 2.3 Proof of Regret Bound
According to Proposition 2, the demand estimation error is relatively small with high probability. To explain the bounds in Proposition 2, we note that each bound consists of two terms: The first term is due to the gap between the empirically estimated probability and the underlying true probability; the second term is due to the deviation of the random future demand from the empirically estimated future demand. Note that since the empirical probability is also infrequently updated, the demand estimation $\boldsymbol{d}^{t}$ contains $\hat{\boldsymbol{p}}^{T_{k-1}}$ rather than $\hat{\boldsymbol{p}}^{t}$, and hence the concentration bounds in Proposition 2 contains $T_{k-1}$.
Step 3: Surrogate LP for $\boldsymbol{u}^{t}$. In the literature on argmax policies (e.g., Vera and Banerjee 2020 and Vera et al. 2021), the proposed policy makes decisions based on the optimal solution of the fluid problem at each period. This approach allows them to bridge the proposed policy with the optimal policy using two LPs: the fluid problem and the hindsight problem. However, we solve the fluid problem only at a few selected periods, and approximate the optimal solution by $\boldsymbol{u}^{t}$ in other periods. In this case, $\boldsymbol{u}^{t}$ is not an optimal solution to the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ for most time periods. Consequently, the linkage between the fluid model and the proposed policy breaks down, and the proof technique in the literature cannot directly apply to our proof. To overcome this challenge, we prove that $\boldsymbol{u}^{t}$ is a good approximation to the optimal solutions of the fluid problem by introducing a surrogate LP. In the following, we show that $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP with high probability.
@@
The first part in Proposition 3 is proved by concentration equalities. The second part is proved by induction: At the resolving period $T_{k-1}$, $\boldsymbol{u}^{T_{k-1}}$ is certainly the optimal solution of $\phi(\boldsymbol{b}^{T_{k-1}},\boldsymbol{d}^{T_{k-1}})$. Then, due to the argmax policy and the designed first-order computations in Algorithm 1, we can show that the statement holds for any period $\ell\in(T_{k-1},t]$ as long as $\boldsymbol{d}^{t}\geq\boldsymbol{2}$. Proposition 3 shows, the approximated solution $\boldsymbol{u}^{t}$ is optimal to a surrogate LP, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, with high probability. Recall that we prove that the demand estimation $\boldsymbol{d}^{t}$ is not far away from the random demand $\boldsymbol{Z}^{t}$. Then, we will prove that the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ is not far away from the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$. Thus, we can show that if $u_{j}^{t}\geq d_{j}^{t}/2$, then there exists an optimal solution $\boldsymbol{y}^{*}$ of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$ such that $y^{*}_{j}\geq 1$ with high probability, implying that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j,1)=0$ according to Proposition 1. The logic for the case with $u_{j}^{t}<d_{j}^{t}/2$ is similar.
@@
The proof idea of Proposition 4 is as follows: Consider a type-$j$ customer is accepted at period $t$ and the good event happens. In this case, we have $\boldsymbol{d}^{t}\approx\boldsymbol{Z}^{t}$, $\boldsymbol{d}^{t}\geq\frac{T-t+1}{2}\boldsymbol{p}$ and that $\boldsymbol{u}^{t}$ is an optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$. Since $\boldsymbol{d}^{t}\approx\boldsymbol{Z}^{t}$, we have $\boldsymbol{u}^{t}\approx\tilde{\boldsymbol{y}}^{*}$, where $\tilde{\boldsymbol{y}}^{*}$ denotes the optimal solution to $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$. Then, when $T-t+1$ is greater than a constant, we have $\tilde{y}^{*}_{j}\approx u_{j}^{t}\geq d_{j}^{t}/2\geq\frac{T-t+1}{4}p_{j}\geq 1$, resulting in the results in Proposition 4. According to Proposition 4, for most periods, under the good event, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j^{t}}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$, resulting in $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$. Therefore, the probability $\mathbb{P}\left(\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})>0\right)$ is upper bounded by the bad event probability. Finally, with the above four steps, we are prepared to prove Theorem 1 by bounding the right-hand side of (6).